# More on soundness in the enriched context

### v2

@@ 5. Soundness and universal algebra
In this section we extend the work of [40] to the setting of weakly sound classes of weights. Indeed, while in [40] the aim was to give a notion of $\lambda$-ary language and $\lambda$-ary equational theory whose models characterize the ${\mathcal{V}}$-category of algebras of $\lambda$-ary monads on ${\mathcal{V}}$; here we extend these notions to obtain a characterization of the ${\mathcal{V}}$-category of algebras of those monads on ${\mathcal{V}}$ which preserve $\Phi$-flat colimits, for a given weakly sound class $\Phi$ of weights.
We start by recalling the notion of languages, structures, and terms introduced in [40].
@@
The ${\mathcal{V}}$-category $\operatorname{\textnormal{S}tr}({\mathbb{L}})$ of ${\mathbb{L}}$-structured is defined in [40, Definition 3.3]; it has ${\mathbb{L}}$-structures as objects and morphisms of such as arrows. This comes together with a forgetful ${\mathcal{V}}$-functor $U\colon\operatorname{\textnormal{S}tr}({\mathbb{L}})\to{\mathcal{V}}$ which has a left adjoint $F$.
@@
Given a (possibly extended) term $t:(X,Y)$ and $A\in\operatorname{\textnormal{S}tr}({\mathbb{L}})$, we define the interpretation $t_{A}\colon A^{X}\to A^{Y}$ of $t$ in $A$ following [40, Section 4]. Then one can talk about equations:
Definition 5.4 ([40]).
@@
[…]
@@
${\mathcal{K}}_{0}(A,B^{X})$$\operatorname{\textnormal{M}od}({\mathbb{E}})_{0}(GA,GB^{X})$${\mathcal{V}}_{0}(UA,UB^{X})$$G_{AB^{X}}$$(U_{0})_{AB^{X}}$$\ \ ((U_{\mathbb{E}})_{0})_{AB^{X}}$
@@
${\mathcal{V}}_{0}(X,{\mathcal{K}}(A,B))$${\mathcal{V}}_{0}(X,\operatorname{\textnormal{M}od}({\mathbb{E}})(GA,GB))$${\mathcal{V}}_{0}(X,[UA,UB])$$G^{\prime}_{AB}$${\mathcal{V}}_{0}(X,U_{AB})$$\ \ {\mathcal{V}}_{0}(X,(U_{\mathbb{E}})_{AB})$
@@ 5.1. Languages and theories with respect to a weakly sound class
For the results of the paper it would be enough to assume that ${\mathcal{V}}_{0}$ is locally bounded [24, Section 6] and that has a small dense generator. Essentially by definition, every ${\mathcal{V}}$ which is locally presentable satisfies these two properties. However, whether there are non locally presentable examples (even satisfying just the density hypothesis), is independent from ZFC, and actually equivalent to the Vopenka’s principle [4, 6.B]. Thus, we do not really lose much generality by assuming local presentability, and we can also use the results of [40].
We fix as subcategory of arities $\Phi{\mathcal{I}}$, the closure of =${\mathcal{I}}=\{I\}$ under $\Phi$-colimits in ${\mathcal{V}}$. (The notation is that of the free cocompletion under $\Phi$-colimits).
@@
$[\Phi{\mathcal{I}},{\mathcal{V}}]\xrightarrow{\ \simeq\ }[{\mathcal{V}}_{\Phi},{\mathcal{V}}]$
@@
The following generalizes [40, 5.11] from the $\lambda$-ary to the sound case.
@@
The fact that $U\colon\operatorname{\textnormal{M}od}({\mathbb{E}})\to{\mathcal{V}}$ is strictly monadic follows from [40, 5.11] since every $\Phi$-ary theory is $\lambda$-ary for some $\lambda$. We are left to prove preservation of $\Phi$-flat colimits.
First, let us prove that the forgetful ${\mathcal{V}}$-functor $V\colon\operatorname{\textnormal{S}tr}({\mathbb{L}})\to{\mathcal{V}}$ preserves these. Let $M\colon{\mathcal{C}}^{\textnormal{op}}\to{\mathcal{V}}$ be any $\Phi$-flat weight and $H\colon{\mathcal{C}}\to\operatorname{\textnormal{S}tr}({\mathbb{L}})$; we prove that $M*VH$ has a naturally induced notion of ${\mathbb{L}}$-structure. Indeed, for any function symbol $f:(X,Y)$ in ${\mathbb{L}}$ we can consider the composite
[…]
where we used that $X$ is $\Phi$-presentable in the initial isomorphism, considered the ${\mathcal{V}}$-functor $V_{f}\colon\operatorname{\textnormal{S}tr}({\mathbb{L}})\to{\mathcal{V}}^{\mathbbm{2}}$ evaluating at $f\in{\mathbb{L}}$ in the middle map, and took the comparison morphism as our final map. It now easily follows that $V$ creates $\Phi$-flat colimits, and hence it preserves them.
Consider now the inclusion $J\colon\operatorname{\textnormal{M}od}({\mathbb{E}})\hookrightarrow\operatorname{\textnormal{S}tr}({\mathbb{L}})$. To conclude we need to prove that this also preserves $\Phi$-flat colimits; then also $U$ does since it is the composite of $J$ and $V$. Given any $\Phi$-flat weight $M\colon{\mathcal{C}}^{\textnormal{op}}\to{\mathcal{V}}$ and $H\colon{\mathcal{C}}\to\operatorname{\textnormal{M}od}({\mathbb{E}})$, it is enough to prove that the ${\mathbb{L}}$-structure $M*JH$ still satisfies the equations of ${\mathbb{E}}$. Take any $(s=t)$ in ${\mathbb{E}}$, where $s$ and $t$ are $(X,Y)$-ary with $X\in\Phi{\mathcal{I}}$ by hypothesis; then we can consider the square
[…]
where we are using that the underlying object of $M*JH$ is $M*UH$ (this follows from the first part of the proof), that $q$ and $q^{\prime}$ are comparison maps and $q$ is an isomorphism since $X\in\Phi{\mathcal{I}}$, and $U_{s},U_{t}\colon\operatorname{\textnormal{S}tr}({\mathbb{L}})\to{\mathcal{V}}^{\mathbbm{2}}$ are evaluating at $s$ and $t$. Now, since $H$ lands in $\operatorname{\textnormal{M}od}({\mathbb{E}})$, it follows that $(s_{HC}=t_{HC})$ for any $C\in{\mathcal{C}}$, or equivalently (in the notation above) that $U_{s}HC=U_{t}HC$ for any $C$. Therefore we also obtain the equality $M*U_{s}H=M*U_{t}H$; since $q$ is an isomorphism this implies that $M*JH$ satisfies $(s=t)$. ∎
For other implication we can apply the results of [15, 35] to $\Phi{\mathcal{I}}$ since a ${\mathcal{V}}$-functor ${\mathcal{V}}\to{\mathcal{V}}$ is the left Kan extension of its restriction to $\Phi{\mathcal{I}}$ if and only if it preserves $\Phi$-flat colimits (Proposition 5.8).
@@
Let $T\colon{\mathcal{V}}\to{\mathcal{V}}$ be a monad preserving $\Phi$-flat colimits. Then there exists a $\Phi$-ary equational theory ${\mathbb{E}}$ on a $\Phi$-ary language ${\mathbb{L}}$ together with an isomorphism $E\colon\textnormal{Alg}(T)\to\operatorname{\textnormal{M}od}({\mathbb{E}})$ making the triangle
@@
The proof is somewhat inspired by that of [40, 5.12] but requires some subtle modifications since we cannot use [30, 9.2] because $(\Phi{\mathcal{I}})_{0}$ is not a strong generator of ${\mathcal{V}}_{0}$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 5.2. Characterization theorem
In this section we use the results above to prove the main characterization theorem for ${\mathcal{V}}$-categories of models of equational $\Phi$-theories. Recall from [40] that an object $G$ of a ${\mathcal{V}}$-category ${\mathcal{K}}$ is called ${\mathcal{V}}$-projective if ${\mathcal{K}}(G,-)$ preserves coequalizers of ${\mathcal{K}}(G,-)$-split pairs.
@@
[…]
@@
Same argument used to prove [40, Proposition 5.19], with the only difference that now $X\in\Phi{\mathcal{I}}$. ∎
@@ Appendix A Sifted colimits for ${\mathcal{V}}=\mathbf{Cat}$
Let ${\mathcal{C}}$ be a 2-category with finite coproducts; then the free cocompletion $\mathbf{Ref}({\mathcal{C}})$ is finitely cocomplete and the inclusion $J\colon{\mathcal{C}}\hookrightarrow\mathbf{Ref}({\mathcal{C}})$ preserves finite coproducts. Moreover, if ${\mathcal{A}}$ is finitely cocomplete, a 2-functor $M\colon{\mathcal{C}}\to{\mathcal{A}}$ preserves finite coproducts if and only if $\textnormal{Lan}_{J}M$ is finitely cocontinuous.
@@
It is easy to see that $\mathbf{Ref}({\mathcal{C}})$ can be defined as follows: consider ${\mathcal{D}}_{0}={\mathcal{C}}$, and for $k\geq 0$ define ${\mathcal{D}}_{2k+1}$ as the full subcategory of $[{\mathcal{C}}^{\textnormal{op}},\mathbf{Cat}]$ spanned by the reflexive coequalizers of elements of ${\mathcal{D}}_{2k}$, and ${\mathcal{D}}_{2k+2}$ as the full subcategory of $[{\mathcal{C}}^{\textnormal{op}},\mathbf{Cat}]$ spanned by the reflexive codescent objects of elements of ${\mathcal{D}}_{2k+1}$. Then $\mathbf{Ref}({\mathcal{C}})=\cup_{n}{\mathcal{D}}_{n}$.
By Lemma A.2 above, to prove that $\mathbf{Ref}({\mathcal{C}})$ is finitely complete it is enough to show that it has finite coproducts. Since $\mathbf{Ref}$-colimits are 2-sifted it follows that $\mathbf{Ref}({\mathcal{C}})\subseteq\mathbf{2}\textnormal{-}\mathbf{Sind}({\mathcal{C}})$; thus it is enough to show that $\mathbf{Ref}({\mathcal{C}})$ is closed in $\mathbf{2}\textnormal{-}\mathbf{Sind}({\mathcal{C}})$ under finite coproducts (which exist in $\mathbf{2}\textnormal{-}\mathbf{Sind}({\mathcal{C}})$ being cocomplete). We do that by showing that ${\mathcal{D}}_{n}$ is closed under finite coproducts in $\mathbf{2}\textnormal{-}\mathbf{Sind}({\mathcal{C}})$ for each $n\in\mathbb{N}$.
For $n=0$ the result is true simply because ${\mathcal{D}}_{0}={\mathcal{C}}$ has finite coproducts and 2-sifted colimits commute in $\mathbf{Cat}$ with finite products. Consider now $n=2k+1$ and assume by inductive hypothesis that ${\mathcal{D}}_{2k}$ is closed under finite coproducts in $\mathbf{2}\textnormal{-}\mathbf{Sind}({\mathcal{C}})$. Given $X,Y\in{\mathcal{D}}_{2k+1}$, we can write $X$ and $Y$ as reflexive coequalizers of maps in ${\mathcal{D}}_{2k}$; take the coproduct of these maps in ${\mathcal{D}}_{2k}$, then their coequalizer $Z$ in $\mathbf{2}\textnormal{-}\mathbf{Sind}({\mathcal{C}})$ is the coproduct of $X$ and $Y$ (by the commutativity of colimits with other colimits and the fact that ${\mathcal{D}}_{n}$ is closed under finite coproducts). Since reflexive coequalizers in ${\mathcal{D}}_{2k+1}$ are computed as in $\mathbf{2}\textnormal{-}\mathbf{Sind}({\mathcal{C}})$, it follows that $Z$ actually lies in ${\mathcal{D}}_{2k+1}$ and is the coproduct of $X$ and $Y$. The same argument applies for ${\mathcal{D}}_{2k+2}$.
As a consequence $\mathbf{Ref}({\mathcal{C}})$ is finitely cocomplete and the inclusion $J\colon{\mathcal{C}}\hookrightarrow\mathbf{Ref}({\mathcal{C}})$ preserves finite coproducts. Moreover, by the explicit construction of coproducts given above, it follows that if a 2-functor $M\colon{\mathcal{C}}\to{\mathcal{A}}$ preserves finite coproducts then $\textnormal{Lan}_{J}M$ preserves them too. The last part of the statement is then a consequence of Lemma A.2 and the universal property of $\mathbf{Ref}$-cocompletions. ∎
@@
Let ${\mathcal{C}}$ be any 2-category; then $\mathbf{Fin}({\mathcal{C}})\simeq\mathbf{Ref}(\mathbf{Fp}({\mathcal{C}}))$.
@@
$Z\colon{\mathcal{C}}\hookrightarrow\mathbf{Fp}({\mathcal{C}})\hookrightarrow\mathbf{Ref}(\mathbf{Fp}({\mathcal{C}})),$
@@
Below we denote by $\mathbf{Ind}(-)$ the free cocompletion under filtered colimits.
@@
$\mathbf{2}\textnormal{-}\mathbf{Sind}({\mathcal{C}})\simeq\mathbf{Ind}(\mathbf{Ref}({\mathcal{C}})).$
The equivalence being induced by the map $\mathbf{Ind}(\mathbf{Ref}({\mathcal{C}}))\to\mathbf{2}\textnormal{-}\mathbf{Sind}({\mathcal{C}})$ obtained by left Kan extending the inclusion ${\mathcal{C}}\hookrightarrow\mathbf{2}\textnormal{-}\mathbf{Sind}({\mathcal{C}})$.
@@
$\begin{split}\mathbf{2}\textnormal{-}\mathbf{Sind}({\mathcal{C}})&\simeq\mathbf{Fp}({\mathcal{C}}^{\textnormal{op}},\mathbf{Cat})\\ &\simeq\mathbf{Lex}(\mathbf{Ref}({\mathcal{C}})^{\textnormal{op}},\mathbf{Cat})\\ &\simeq\mathbf{Ind}(\mathbf{Ref}({\mathcal{C}}))\\ \end{split}$
where the first equivalence is given by soundness of the class of finite products, the second is obtained by right Kan extending along $J^{\textnormal{op}}\colon{\mathcal{C}}^{\textnormal{op}}\hookrightarrow\mathbf{Ref}({\mathcal{C}})^{\textnormal{op}}$ and is a consequence of (the dual of) Proposition A.4. The third equivalence is given by soundness of the class of finite limits, plus the fact that freely adding filtered colimits is the same as adding $\mathbf{Lex}$-flat colimits (by [31, Theorem 3.13] applied to ${\mathcal{V}}=\mathbf{Cat}$). ∎
@@
Let ${\mathcal{K}}$ be cocomplete and consider a sifted weight $M\colon{\mathcal{C}}^{\textnormal{op}}\to\mathbf{Cat}$ together with a diagram $H\colon{\mathcal{C}}\to{\mathcal{K}}$. Without loss of generality we can assume that ${\mathcal{C}}$ has finite coproducts; indeed, we can replace ${\mathcal{C}}$ with its free cocompletion $\mathbf{Fp}({\mathcal{C}})$ under finite coproducts, $M$ with $M^{\prime}:=\textnormal{Lan}_{J^{\textnormal{op}}}M$ and $H$ with $H^{\prime}:=\textnormal{Lan}_{J}H$, where $J\colon{\mathcal{C}}\to\mathbf{Fp}({\mathcal{C}})$ is the inclusion (then $M^{\prime}$ is still 2-sifted and $M*H\cong M^{\prime}*H^{\prime}$).
Now, it follows that $M\in\mathbf{2}\textnormal{-}\mathbf{Sind}({\mathcal{C}})$, which by Corollary A.6 above is equivalent to $\mathbf{Ind}(\mathbf{Ref}({\mathcal{C}}))$. Therefore $M$ is a filtered colimit of $\mathbf{Ref}$-colimits of representables. Thus, the preservation of $M$-weighted colimits follows from that of filtered colimits and $\mathbf{Ref}$-colimits, as claimed.

### v3

@@ 5. Soundness and universal algebra
In this section we extend the work of [42] to the setting of weakly sound classes of weights. Indeed, while in [42] the aim was to give a notion of $\lambda$-ary language and $\lambda$-ary equational theory (based on recursively generated terms) whose models characterize the ${\mathcal{V}}$-category of algebras of $\lambda$-ary monads on ${\mathcal{V}}$; here we extend these notions to obtain a characterization of the ${\mathcal{V}}$-category of algebras of those monads on ${\mathcal{V}}$ which preserve $\Phi$-flat colimits, for a given weakly sound class $\Phi$ of weights.
We start by recalling the notion of languages, structures, and terms introduced in [42]. The same notions of languages and structures were considered before in [36, Section 5].
@@
The ${\mathcal{V}}$-category $\operatorname{\textnormal{S}tr}({\mathbb{L}})$ of ${\mathbb{L}}$-structured is defined in [42, Definition 3.3]; it has ${\mathbb{L}}$-structures as objects and morphisms of such as arrows. This comes together with a forgetful ${\mathcal{V}}$-functor $U\colon\operatorname{\textnormal{S}tr}({\mathbb{L}})\to{\mathcal{V}}$ which has a left adjoint $F$.
The recursively generated terms generalize those introduced in [42, Definition 4.1], while extended terms were already introduced in [36, Definition 3.2] (see [42, Remark 4.6]).
@@
Given a (possibly extended) term $t:(X,Y)$ and $A\in\operatorname{\textnormal{S}tr}({\mathbb{L}})$, we define the interpretation $t_{A}\colon A^{X}\to A^{Y}$ of $t$ in $A$ following [42, Section 4]. Then one can talk about equations:
Definition 5.4 ([36, 42]).
@@
${\mathcal{K}}_{0}$$\operatorname{\textnormal{M}od}({\mathbb{E}})_{0}$${\mathcal{V}}$$G$$U_{0}$$(U_{\mathbb{E}})_{0}$
@@
${\mathcal{K}}_{0}(A,B^{X})$$\operatorname{\textnormal{M}od}({\mathbb{E}})_{0}(GA,GB^{X})$${\mathcal{V}}_{0}(UA,UB^{X})$$G_{AB^{X}}$$(U_{0})_{AB^{X}}\ \$$\ \ ((U_{\mathbb{E}})_{0})_{AB^{X}}$
@@
${\mathcal{V}}_{0}(X,{\mathcal{K}}(A,B))$${\mathcal{V}}_{0}(X,\operatorname{\textnormal{M}od}({\mathbb{E}})(GA,GB))$${\mathcal{V}}_{0}(X,[UA,UB])$$G^{\prime}_{AB}$${\mathcal{V}}_{0}(X,U_{AB})\ \$$\ \ {\mathcal{V}}_{0}(X,(U_{\mathbb{E}})_{AB})$
@@ 5.1. Languages and theories with respect to a weakly sound class
For the results of the paper it would be enough to assume that ${\mathcal{V}}_{0}$ is locally bounded [25, Section 6] and that has a small dense generator. Essentially by definition, every ${\mathcal{V}}$ which is locally presentable satisfies these two properties. However, whether there are non locally presentable examples (even satisfying just the density hypothesis), is independent from ZFC, and actually equivalent to the Vopenka’s principle [4, 6.B]. Thus, we do not really lose much generality by assuming local presentability, and we can also use the results of [42].
We fix as subcategory of arities $\Phi{\mathcal{I}}$, the closure of ${\mathcal{I}}=\{I\}$ under $\Phi$-colimits in ${\mathcal{V}}$. (The notation is that of the free cocompletion under $\Phi$-colimits).
@@
$\Phi{\mathcal{I}},{\mathcal{V}}]\xrightarrow{\ \simeq\ }[{\mathcal{V}}_{\Phi},{\mathcal{V}}]$
@@
The following generalizes [42, 5.11] from the $\lambda$-ary to the sound case, and is a consequence of [36, Theorem 5.20]
@@
Given [36, Example 2.4(5)], this follows from [36, Theorem 5.20]. ∎
For other implication we cannot use the results of [36], as we want to characterize ${\mathcal{V}}$-categories of algebras as models of equational theories which contain only standard (that is, recursively generated) terms. Nonetheless, we can apply the results of [15, 37] to $\Phi{\mathcal{I}}$ since a ${\mathcal{V}}$-functor ${\mathcal{V}}\to{\mathcal{V}}$ is the left Kan extension of its restriction to $\Phi{\mathcal{I}}$ if and only if it preserves $\Phi$-flat colimits (Proposition 5.8).
@@
Let $T\colon{\mathcal{V}}\to{\mathcal{V}}$ be a monad preserving $\Phi$-flat colimits. Then there exists a $\Phi$-ary equational theory ${\mathbb{E}}$, whose equations are defined between standard terms, on a $\Phi$-ary language ${\mathbb{L}}$ together with an isomorphism $E\colon\textnormal{Alg}(T)\to\operatorname{\textnormal{M}od}({\mathbb{E}})$ making the triangle
@@
The proof is somewhat inspired by that of [42, 5.12] but requires some subtle modifications since we cannot use [31, 9.2] because $(\Phi{\mathcal{I}})_{0}$ is not a strong generator of ${\mathcal{V}}_{0}$.
@@
$\textnormal{Alg}(T)\displaystyle\lrcorner$$[{\mathcal{T}},{\mathcal{V}}]$${\mathcal{V}}$$[\Phi{\mathcal{I}}^{\textnormal{op}},{\mathcal{V}}]$$U$$[H,{\mathcal{V}}]$${\mathcal{V}}(K-,1)$
@@
$\textnormal{Alg}(T)\displaystyle\lrcorner$$\Phi{\mathcal{I}}\textnormal{-Pw}[{\mathcal{T}},{\mathcal{V}}]$${\mathcal{V}}$$\Phi{\mathcal{I}}\textnormal{-Pw}[\Phi{\mathcal{I}}^{\textnormal{op}},{\mathcal{V}}]$$U$$[H,{\mathcal{V}}]$${\mathcal{V}}(K-,1)$
@@
${\mathcal{T}}(X_{2},X_{3})\otimes{\mathcal{T}}(X_{1},X_{2})$${\mathcal{T}}(X_{1},X_{3})$$[A^{X_{2}},A^{X_{3}}]\otimes[A^{X_{1}},A^{X_{2}}]$$[A^{X_{1}},A^{X_{3}}]$$M$$F_{X_{2},X_{3}}\otimes F_{X_{1},X_{2}}$$F_{X_{1},X_{3}}$$M$
@@
$\operatorname{\textnormal{M}od}(\mathbb{E})_{0}\displaystyle\lrcorner$$\Phi{\mathcal{I}}\textnormal{-Pw}[{\mathcal{T}},{\mathcal{V}}]_{0}$${\mathcal{V}}$$\Phi{\mathcal{I}}\textnormal{-Pw}[\Phi{\mathcal{I}}^{\textnormal{op}},{\mathcal{V}}]_{0}$$(U_{\mathbb{E}})_{0}$$[H,{\mathcal{V}}]_{0}$${\mathcal{V}}(K-,1)_{0}$
@@ 5.2. Characterization theorem
In this section we use the results above to prove the main characterization theorem for ${\mathcal{V}}$-categories of models of equational $\Phi$-theories. Recall from [42] that an object $G$ of a ${\mathcal{V}}$-category ${\mathcal{K}}$ is called ${\mathcal{V}}$-projective if ${\mathcal{K}}(G,-)$ preserves coequalizers of ${\mathcal{K}}(G,-)$-split pairs. Note that $(1)\Leftrightarrow(3)$ was shown as [36, Theorem 5.26]; the improvement here is that it is enough to use the recursively generated terms.
@@
${\mathcal{K}}\displaystyle\lrcorner$$[{\mathcal{T}},{\mathcal{V}}]$${\mathcal{V}}$$[\Phi{\mathcal{I}}^{\textnormal{op}},{\mathcal{V}}]$$U$$[H,{\mathcal{V}}]$${\mathcal{V}}(K-,1)$
@@
Same argument used to prove [42, Proposition 5.19], with the only difference that now $X\in\Phi{\mathcal{I}}$. ∎
@@ Appendix A Sifted colimits for ${\mathcal{V}}=\mathbf{Cat}$
Let ${\mathcal{C}}$ be a 2-category with finite coproducts; then the free cocompletion $\textnormal{Ref}({\mathcal{C}})$ is finitely cocomplete and the inclusion $J\colon{\mathcal{C}}\hookrightarrow\textnormal{Ref}({\mathcal{C}})$ preserves finite coproducts. Moreover, if ${\mathcal{A}}$ is finitely cocomplete, a 2-functor $M\colon{\mathcal{C}}\to{\mathcal{A}}$ preserves finite coproducts if and only if $\textnormal{Lan}_{J}M$ is finitely cocontinuous.
@@
It is easy to see that $\textnormal{Ref}({\mathcal{C}})$ can be defined as follows: consider ${\mathcal{D}}_{0}={\mathcal{C}}$, and for $k\geq 0$ define ${\mathcal{D}}_{2k+1}$ as the full subcategory of $[{\mathcal{C}}^{\textnormal{op}},\mathbf{Cat}]$ spanned by the reflexive coequalizers of elements of ${\mathcal{D}}_{2k}$, and ${\mathcal{D}}_{2k+2}$ as the full subcategory of $[{\mathcal{C}}^{\textnormal{op}},\mathbf{Cat}]$ spanned by the reflexive codescent objects of elements of ${\mathcal{D}}_{2k+1}$. Then $\textnormal{Ref}({\mathcal{C}})=\cup_{n}{\mathcal{D}}_{n}$.
By Lemma A.2 above, to prove that $\textnormal{Ref}({\mathcal{C}})$ is finitely cocomplete it is enough to show that it has finite coproducts. Since $\mathbf{Ref}$-colimits are 2-sifted it follows that $\textnormal{Ref}({\mathcal{C}})\subseteq\textnormal{2}\textnormal{-}\textnormal{Sind}({\mathcal{C}})$; thus it is enough to show that $\textnormal{Ref}({\mathcal{C}})$ is closed in $\textnormal{2}\textnormal{-}\textnormal{Sind}({\mathcal{C}})$ under finite coproducts (which exist in $\textnormal{2}\textnormal{-}\textnormal{Sind}({\mathcal{C}})$ being cocomplete). We do that by showing that ${\mathcal{D}}_{n}$ is closed under finite coproducts in $\textnormal{2}\textnormal{-}\textnormal{Sind}({\mathcal{C}})$ for each $n\in\mathbb{N}$.
For $n=0$ the result is true simply because ${\mathcal{D}}_{0}={\mathcal{C}}$ has finite coproducts and 2-sifted colimits commute in $\mathbf{Cat}$ with finite products. Consider now $n=2k+1$ and assume by inductive hypothesis that ${\mathcal{D}}_{2k}$ is closed under finite coproducts in $\textnormal{2}\textnormal{-}\textnormal{Sind}({\mathcal{C}})$. Given $X,Y\in{\mathcal{D}}_{2k+1}$, we can write $X$ and $Y$ as reflexive coequalizers of maps in ${\mathcal{D}}_{2k}$; take the coproduct of these maps in ${\mathcal{D}}_{2k}$, then their coequalizer $Z$ in $\textnormal{2}\textnormal{-}\textnormal{Sind}({\mathcal{C}})$ is the coproduct of $X$ and $Y$ (by the commutativity of colimits with other colimits and the fact that ${\mathcal{D}}_{n}$ is closed under finite coproducts). Since reflexive coequalizers in ${\mathcal{D}}_{2k+1}$ are computed as in $\textnormal{2}\textnormal{-}\textnormal{Sind}({\mathcal{C}})$, it follows that $Z$ actually lies in ${\mathcal{D}}_{2k+1}$ and is the coproduct of $X$ and $Y$. The same argument applies for ${\mathcal{D}}_{2k+2}$.
As a consequence $\textnormal{Ref}({\mathcal{C}})$ is finitely cocomplete and the inclusion $J\colon{\mathcal{C}}\hookrightarrow\textnormal{Ref}({\mathcal{C}})$ preserves finite coproducts. Moreover, by the explicit construction of coproducts given above, it follows that if a 2-functor $M\colon{\mathcal{C}}\to{\mathcal{A}}$ preserves finite coproducts then $\textnormal{Lan}_{J}M$ preserves them too. The last part of the statement is then a consequence of Lemma A.2 and the universal property of $\mathbf{Ref}$-cocompletions. ∎
@@
Let ${\mathcal{C}}$ be any 2-category; then $\textnormal{Fin}({\mathcal{C}})\simeq\textnormal{Ref}(\textnormal{Fp}({\mathcal{C}}))$.
@@
$Z\colon{\mathcal{C}}\hookrightarrow\textnormal{Fp}({\mathcal{C}})\hookrightarrow\textnormal{Ref}(\textnormal{Fp}({\mathcal{C}})),$
@@
Below we denote by $\textnormal{Ind}(-)$ the free cocompletion under filtered colimits.
@@
$\textnormal{2}\textnormal{-}\textnormal{Sind}({\mathcal{C}})\simeq\textnormal{Ind}(\textnormal{Ref}({\mathcal{C}})).$
The equivalence being induced by the map $\textnormal{Ind}(\textnormal{Ref}({\mathcal{C}}))\to\textnormal{2-Sind}({\mathcal{C}})$ obtained by left Kan extending the inclusion ${\mathcal{C}}\hookrightarrow\textnormal{2-Sind}({\mathcal{C}})$.
@@
[…]
where the first equivalence is given by soundness of the class of finite products, the second is obtained by right Kan extending along $J^{\textnormal{op}}\colon{\mathcal{C}}^{\textnormal{op}}\hookrightarrow\textnormal{Ref}({\mathcal{C}})^{\textnormal{op}}$ and is a consequence of (the dual of) Proposition A.4. The third equivalence is given by soundness of the class of finite limits, plus the fact that freely adding filtered colimits is the same as adding $\mathbf{Lex}$-flat colimits (by [32, Theorem 3.13] applied to ${\mathcal{V}}=\mathbf{Cat}$). ∎
@@
Let ${\mathcal{K}}$ be cocomplete and consider a sifted weight $M\colon{\mathcal{C}}^{\textnormal{op}}\to\mathbf{Cat}$ together with a diagram $H\colon{\mathcal{C}}\to{\mathcal{K}}$. Without loss of generality we can assume that ${\mathcal{C}}$ has finite coproducts; indeed, we can replace ${\mathcal{C}}$ with its free cocompletion $\textnormal{Fp}({\mathcal{C}})$ under finite coproducts, $M$ with $M^{\prime}:=\textnormal{Lan}_{J^{\textnormal{op}}}M$ and $H$ with $H^{\prime}:=\textnormal{Lan}_{J}H$, where $J\colon{\mathcal{C}}\to\textnormal{Fp}({\mathcal{C}})$ is the inclusion (then $M^{\prime}$ is still 2-sifted and $M*H\cong M^{\prime}*H^{\prime}$).
Now, it follows that $M\in\textnormal{2-Sind}({\mathcal{C}})$, which by Corollary A.6 above is equivalent to $\textnormal{Ind}(\textnormal{Ref}({\mathcal{C}}))$. Therefore $M$ is a filtered colimit of $\mathbf{Ref}$-colimits of representables. Thus, the preservation of $M$-weighted colimits follows from that of filtered colimits and $\mathbf{Ref}$-colimits, as claimed.